OpenStudy (helpjebal):

tan (arcsin (2/5)) find the exact value. Not sure how to go about this... can anyone point me in the right direction?

zepdrix (zepdrix):

\[\rm \tan\left[\color{orangered}{\arcsin\left(\frac{2}{5}\right)}\right]\]Let's examine the inside first.\[\rm \color{orangered}{\arcsin\left(\frac{2}{5}\right)}\]

zepdrix (zepdrix):

Recall that when you take the `sine` of an `angle`, you end up with the `ratio` of a side to another side, namely, opposite to hypotenuse. So when we take the `inverse sine` of a `ratio`, you end up with an `angle`.\[\rm \color{orangered}{\arcsin\left(\frac{2}{5}\right)=\theta}\]

zepdrix (zepdrix):

We can rewrite this in terms of the sine function, instead of the inverse. Do you remember how to do that?

OpenStudy (helpjebal):

no... is that when I'd need he special calculator?

OpenStudy (helpjebal):

*the

zepdrix (zepdrix):

Here is one approach: Applying sine function to each side gives us:\[\large\rm \sin\left(\arcsin\left(\frac{2}{5}\right)\right)=\sin(\theta)\]Since the left side is the composition of a function and it's inverse, we simply get the argument as a result,\[\large\rm \frac{2}{5}=\sin(\theta)\] If that process is confusing, a simpler way to think about it is... When switching from inverse sine to sine, the two things switch places.\[\rm \color{orangered}{\arcsin\left(\frac{2}{5}\right)=\theta}\qquad\to\qquad \sin (\theta)=\frac{2}{5}\]

zepdrix (zepdrix):

\[\rm \sin(\theta)=\frac{2}{5}=\frac{opposite}{hypotenuse}\]We want to draw a triangle to show this relationship.

zepdrix (zepdrix):

We'll arbitrarily place theta in the bottom left corner here. Do you understand where the 2 and 5 would go in this diagram?

OpenStudy (helpjebal):

the 2 would be on the right, vertical straight part of the triangle and the 5 would be on the diagonal part, right?

zepdrix (zepdrix):

Yes, good. From there, we'll apply our Pythagorean Theorem to find the missing side length. So what do you get for that length? :) No decimal, leave it as an ugly number if that's what you end up with.